Three student teams collected 65 books for the school library. The first team collected 10 fewer books than the second team and the third team collected 30% of the number of books collected by first and second teams together. How many books did each team collect?
Answer by Mini-(17)

Let: X + Y + Z = 65
X = Y - 10
Z = 0.30 * ( X + Y )
Then:
Z = 0.3 * ( (Y - 10) + Y )
Z = 0.3 * ( 2*Y - 10 ) = 0.6*Y - 3
(Y - 10) + Y + 0.6*Y - 3 = 65
2.6*Y - 13 = 65
2.6*Y = 78
Y = 78 / 2.6 = 30
X = Y - 10 = 20
Z = 0.3 * 50 = 15